OpenStudy (anonymous):

Look at the right triangle ABC. A student made the following chart to prove that AB2 + BC2 = AC2.

OpenStudy (anonymous):

Statement Justification 1. Triangle ABC is similar to triangle BDC 1. Angle ABC = Angle BCD and Angle BCA = Angle DBC 2. BC2 = AC × DC 2. BC ÷ DC = AC ÷ BC because triangle ABC is similar to triangle BDC 3. Triangle ABC is similar to triangle ABD 3. Angle ABC = Angle ADB and Angle BAC = Angle BAD 4. AB2 = AC × AD 4. AB ÷ AD = AC ÷ AB because triangle ABC is similar to triangle ABD 5. AB2 + BC2 = AC × AD + AC × DC = AC (AD + DC) 5. Adding Statement 1 and Statement 2 6. AB2 + BC2 = AC2 6. AD + DC = AC What is the flaw in the student’s proof? Justification 4 should be “AB ÷ AD = AB ÷ AC because triangle ABC is similar to triangle ABD”. Justification 1 should be “Angle ABC = Angle BDC and Angle BCA = Angle BCD”. Justification 2 should be “BC ÷ DC = BC ÷ AC because triangle ABC is similar to triangle BDC”. Justification 3 should be “Angle ABC = Angle BAD and Angle BAC = Angle ABD”.

OpenStudy (anonymous):

Is it option #3?

OpenStudy (anonymous):

May someone please help me?

OpenStudy (anonymous):

Angle ABC = Angle BDC* What is BC2? Is that BC^2?

OpenStudy (anonymous):

Yes, it is the latter... BC^2. Sorry. I guess it was a typo.

OpenStudy (anonymous):

No worries. And as you can see the first statement in justification is fixed. That was a flaw because the two 90 degree angles are similar and not a 90 degree and a 45 degree angle.

OpenStudy (anonymous):

Ok. Thanks!!!

OpenStudy (anonymous):

Are there any units? I can't make assumptions for justification 2.

OpenStudy (anonymous):

No. That's all they provided.

OpenStudy (anonymous):

Well we can assume that length BC is one less than AC. Or is AC two units longer?

OpenStudy (anonymous):

The look equal, but I wouldn't assume that. They don't often draw things to scale...

OpenStudy (anonymous):

But I'd say they are equal, sine it is the angle bisector there, and the median of AC

OpenStudy (anonymous):

We can use cosine law to find the lengths of BC and DC.

OpenStudy (phi):

just for clarification 2. BC ÷ DC = AC ÷ BC because triangle ABC is similar to triangle BDC means length(BD) is to length(DC) as length(AC) is to length(BC) i.e. ratios based on corresponding legs of similar triangles.

OpenStudy (anonymous):

Yes. But we can't assume any lengths till we find the exact number which we can manipulate to justify the statement made regarding the similar triangles.

OpenStudy (anonymous):

ok, So, we find the lengths, then do as phi suggested?

OpenStudy (anonymous):

Yes we should. Let me get to a pleasant environment. I hate being near controversies depicted on news channels.

OpenStudy (anonymous):

lol soo random aha.

OpenStudy (anonymous):

Yes... me too. Always one party making unbelievable accusations against the other.... but then again, it's important, because it'll affect the people in this nation, and in other nations(maybe) as well.

OpenStudy (anonymous):

Now, back to math.....

OpenStudy (phi):